40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1442     Entry time: Mon Mar 30 12:29:17 2009
Author: Yoichi 
Type: Configuration 
Category:  
Subject: MC1 glitches not seen during the weekend 
The attached is the MC trend for the past 12 hours.
There is no MC1 glitches in the OSEM signals. Moreover, the total amplitude of the drift is smaller than it used to be (now the amplitude is less than 100 but it used to be a few hundreds).
There still is a small step in the OSEM signals at around 6AM this morning, but the amount of the jump is very insignificant.
The cause of the glitch in the TMP1, DRUM1 and APTEMP (LL, UL and UR coils respectively) at 7AM is not known.
Since the MC1 has been behaving OK during the weekend, I removed the probes from the MC1 coil driver board and locked the MC.
Hopefully the replacement of the broken connector fixed the problem, but I'm not sure.
Attachment 1: MC_drift.pdf  45 kB  | Show | Show all
ELOG V3.1.3-